### v2

# Fast Jet Finding in Julia

Abstract
Jet reconstruction remains a critical task in the analysis of data from HEP colliders. We describe in this paper a new, highly performant, Julia package for jet reconstruction, JetReconstruction.jl, which integrates into the growing ecosystem of Julia packages for HEP. With this package users can run sequential reconstruction algorithms for jets. In particular, for LHC events, the Anti-${k}_{\text{T}}$, Cambridge/Aachen and Inclusive-${k}_{\text{T}}$ algorithms can be used. For FCCee studies the use of alternative algorithms such as the Generalised ${k}_{\text{T}}$ for $e^{+}e^{-}$ and Durham are also supported.
The performance of the core algorithms is better than Fastjet’s C++ implementation, for typical LHC and FCCee events, thanks to the Julia compiler’s exploitation of single-instruction-multiple-data (SIMD), as well as ergonomic compact data layouts.
The full reconstruction history is made available, allowing inclusive and exclusive jets to be retrieved. The package also provides the means to visualise the reconstruction. Substructure algorithms have been added that allow advanced analysis techniques to be employed. The package can read event data from EDM4hep files and reconstruct jets from these directly, opening the door to FCCee and other future collider studies in Julia.
1 Introduction
High energy physics (HEP) software is inherently multi-lingual pivarski2022 . Across the field actively used codes exist in many different languages. For the code that is used in mainline HEP workflows two languages have dominated in the last few decades: C++ and Python. C++ saw adoption at the BaBar experiment at SLAC and, subsequently, at the Large Hadron Collider (LHC) experiments. Python has grown enormously in popularity in recent years, becoming almost ubiquitous across the field. These languages have different strengths, with C++ excelling at runtime performance, used heavily for simulation and reconstruction; and Python shining in the areas of rapid turnaround, prototyping and steering, being then particularly strong in the analysis domain.
The current status of C++ and Python, with these languages having rather different paradigms, leads to friction and potentially awkward interfaces. Code developed in Python may not run efficiently at scale, leading to inefficient use of computing resources, even necessitating a rewrite in C++. The current generation of physicists is generally far more comfortable in Python and there is a loss of skills in C++, which is a challenge for the experiments.
An alternative option, which is attracting increasing interest, is to use a language that can bring the runtime advantages of C++, but the ergonomic advantages of Python. The Julia Programming Language bib:julia_freshapproach ; 10.1145/3276490 was designed specifically to do this efficiently and effectively, and has been adopted by many scientific communities perkel-julia-science . In HEP, explorations of Julia have been promising Stanitzki:2020bnx ; eschle2023potential . In particular, a recent comparison of Julia, Python and C++ for the task of sequential jet finding polyglot-jets-chep23 found that Julia performed as well as, or better, than C++, with improved code ergonomics.
In this paper we report on the developments that have happened in the Julia code presented in polyglot-jets-chep23 , in particular the improvements that have resulted in the recent release of the production Julia package, JetReconstruction.jl jetreconstruction-jl-github . The package has been made more accessible to users, with comprehensive documentation and consistent interfaces. New algorithms have been introduced targeting jet reconstruction at $e^{+}e^{-}$ experiments, including reading data from Key4HEP’s EDM4hep Gaede:2022leb data format files. Support for jet substructure analysis at $pp$ colliders has been introduced. We also give the latest benchmarking results, that continue to demonstrate better performance than Fastjet Cacciari:2011ma for almost all parameters.
2 Production Release of Julia Jet Reconstruction
Details of the algorithms and strategies used for $pp$ events in the Julia version of jet finding have been described in polyglot-jets-chep23 . However, due to the different development history of how the two strategies, N2Plain and N2Tiled, were implemented, the original return values from the reconstruction were different. In the N2Plain case an implicit $p_{T}$ cut, selecting inclusive jets, was made; whereas in the N2Tiled strategy a dedicated object that encapsulates the reconstruction sequence was returned that contains a record of all jet merging steps, with parents and metric distance ($d_{ij}$) parameter, allowing the complete graph of the process to be obtained. This structure, a ClusterSequence object, is much more useful as this information is required for further jet processing. This object is now used in all sequential reconstruction algorithms (but always through accessors, to hide internal implementation details).
This choice then allowed the implementation of the other core jet selection, viz. inclusive jets. An interface was added where the same method can be used to make a selection on either the number of final jets, or on the maximum value of the metric distance ($d_{ij}$) – this takes advantage of the fact that in Julia method parameters can be named, providing a clearer interface than the type based method selection in C++ (e.g., inclusive_jets(cs::ClusterSequence; ptmin=5.0) for a $p_{T}$ cut).
In addition to the jet selection, another algorithm for $pp$ jet reconstruction was added, the generalised ${k}_{\text{T}}$ algorithm. This algorithm uses as a momentum metric $k^{2p}_{\text{T}}$, where the power value $p$ is arbitrary (and for specific integer powers of $-1$, 00 and $1$, maps to the well known anti-${k}_{\text{T}}$, Cambridge/Aachen, and inclusive-${k}_{\text{T}}$ algorithms, respectively Cacciari:2005hq; Matteo_Cacciari_2008; fastjetmanual ).
Support for the visualisation of jet reconstruction, through the Julia visualisation package Makie Danisch2021 , has been improved. As Makie is a heavy dependency, we take advantage of the extensions feature of the Julia packaging system, where the visualisation extensions to JetReconstruction.jl are only loaded if Makie already exists in the current user environment. An example of the output from the visualisation extension is shown in Figure 1. Taking advantage of the fact that all reconstruction steps are captured by the ClusterSequence a new visualisation option was added, which animates the reconstruction process jetrecoAnimationCHEP2024 .
Figure 1: Visualisation of a typical $pp$ collision jet reconstruction, using anti-${k}_{\text{T}}$ with $R=2$, in the $y-\phi$ (rapidity, azimuthal angle) plane. The height of each bar indicates the original cluster energy and the colour represents the final jet clustering, i.e., all clusters with the same colour are clustered together.
Of the two core strategies, N2Tiled scales better to higher initial cluster densities, as found in typical LHC $pp$ events. However, there is an overhead for this tiling, which makes the N2Plain strategy better at low cluster densities. It is highly desirable that the user would not have to manually select a strategy, so a heuristic performance scan was made, indicating that the performance of each strategy is about the same for 80 input clusters. So a third strategy, Best was introduced, which selects N2Plain for $<=80$ clusters, otherwise N2Tiled.
Before a useful release of the software could be made, documentation for the package had to be written. This was done using the standard Julia documentation support package, Documenter.jl documenter-jl , which has the great advantage of using the inline code docstrings to document methods. Documentation was then published onto the JuliaHEP organisation’s GitHub Pages website jetreco-docs .
With all of this refactoring done, and with an enhanced suite of tests added, the first public release of JetReconstruction.jl  v0.3.0, was made in June 2024. The package was added to the Julia public registry, making installing it for any user as simple as add JetReconstruction from the standard Julia package interface.
3 Support for $e^{+}e^{-}$ algorithms
To add support for the reconstruction of jets in $e^{+}e^{-}$ events, some different algorithms are needed. The essential idea of sequential jet reconstruction remains the same: calculation of a distance metric between all clusters, then merging the clusters with the lowest metric and repeating. However, for $e^{+}e^{-}$ events it is preferable to reconstruct in geometric space, $(\theta,\phi)$, instead of rapidity space, $(y,\phi)$. This is because experiments usually operate at the production threshold of the processes of particular interest, so jets are less boosted than at the LHC.
There are two main algorithms of interest: the Durham Algorithm and the Generalised ${k}_{\text{T}}$ for $e^{+}e^{-}$.
3.1 Durham Algorithm
For the Durham Algorithm the metric distance between jets $i$ and $j$ is defined as:
$d_{ij}=2\text{min}(E_{i}^{2},E_{j}^{2})(1-\cos\theta_{ij})$
where $E_{i}$ is the cluster energy and $\theta_{ij}$ is the angular separation between $i$ and $j$.
The reconstruction implementation in JetReconstruction.jl is then called with
{minted}
julia cs = jet_reconstruct(particles; algorithm=Durham)
No additional parameters are required. In principle the Durham algorithm proceeds until all clusters are merged to a single jet, but actual analysis will utilise an exclusive jet cut, for the number of jets of interest.
3.2 Generalised ${k}_{\text{T}}$ for $e^{+}e^{-}$
For the Generalised ${k}_{\text{T}}$ for $e^{+}e^{-}$ algorithm, the distance metrics are:
$d_{ij}=\text{min}(E_{i}^{2p},E_{j}^{2p})\frac{1-\cos\theta_{ij}}{1-\cos R}\quad;\quad d_{iB}=E_{i}^{2p}$
For a power value $p$ and a radius parameter $R$. This means jets are finalised when no clusters are found within an angular distance $R$, when $R<\pi$ (this is very similar to the behaviour of the $pp$ algorithms). In JetReconstruction.jl we follow the Fastjet prescription that for $R>\pi$ the denominator is replaced by $3+\cos R$ fastjetmanual .
In the case when $p=1$ and $\pi<R<3\pi$ the clustering sequence is identical to the Durham Algorithm, save for a normalisation factor of 2.
3.3 Implementation Details
For an optimal implementation of the $e^{+}e^{-}$ algorithms we introduce a new Julia structure to represent a jet to be reconstructed in $(\theta,\phi)$ space, an EEjet.
This structure mainly differs from the PseudoJet used for $pp$ reconstruction in that the cached values are optimised for the different reconstruction scheme. Both EEjet and PseudoJet are subtypes of the abstract type FourMomentum, which allows us to parameterise the ClusterSequence on the jet type, thus benefiting from a type specific implementation at runtime, with generic code for the ClusterSequence, viz.
{minted}
julia struct ClusterSequenceT <: FourMomentum … jets::VectorT … end
As the algorithm is being executed additional compact arrays, that track parameters used by the calculation (e.g., the nearest neighbour active cluster, the angle to the nearest neighbour, the fractional momentum in the $x$, $y$ and $z$ directions, etc.), are maintained for computational efficiency. We use the Julia package StructArrays.jl structarrays , which allows us to maintain an ergonomic array of structures interface, but underlying this is a computationally efficient structure of arrays, which gives excellent performance, as shown in Section 4.2.
4 $pp$ and $e^{+}e^{-}$erformance
We have benchmarked the performance of JetReconstruction.jl v0.4.3 against Fastjet v3.4.3111The benchmark machine used was an AMD Ryzen 7, 5700G 3.8GHz (8 cores, plus HT), 32GB RAM, running AlmaLinux 9.4. Julia v1.11.1 was used for JetReconstruction.jl and Fastjet was compiled with gcc 11.4.1 using -O2. Benchmark runs are repeated 32 times and are stable to 1%.. Pythia was used to generate input events at different cluster densities. Each sample consisted of 100 events and where the average density varied from 43 to 632, for $pp$ algorithms and from 43 to 65 for $e^{+}e^{-}$ algorithms.
4.1 $pp$ Algorithms
The results for anti-${k}_{\text{T}}$ jet reconstruction are shown in Figures 2 and 3. The results from JetReconstruction.jl are consistently faster than Fastjet. For the more relevant tiled strategy, which is favoured at $pp$ event densities, Julia is 14-18% faster, depending on the particular algorithm chosen, as seen in Table 1.
Figure 2: Jet reconstruction time for anti-${k}_{\text{T}}$ jet reconstruction at different values of $R$, using the N2Tiled strategy. Reconstruction time for Julia and Fastjet are plotted against the average cluster density of different samples.
Figure 3: Jet reconstruction time for anti-${k}_{\text{T}}$ jet reconstruction at different values of $R$, using the N2Plain strategy. Reconstruction time for Julia and Fastjet are plotted against the average cluster density of different samples.
Algorithm N2Tiled N2Plain N2Plain, $R>=1.0$
Anti-${k}_{\text{T}}$ 1.14 6.46 1.49
Cambridge/Aachen 1.18 6.03 1.80
Inclusive ${k}_{\text{T}}$ 1.16 6.50 1.73
Table 1: Mean ratio of Fastjet reconstruction time to Julia (thus $>1$ indicates Julia is faster) for different $pp$ algorithms, for all values of $R\in(0.2,0.4,1,1.5,2,4)$ and all sample cluster densities. As Fastjet’s behaviour at $R<1$ for N2Plain is pathological, the mean ratio for $R>=1$ is also given.
4.2 $e^{+}e^{-}$ Algorithms
The performance of the Julia versions of the Durham compared to Fastjet is shown in Figures 4(a) and 4(b). The Julia implementation in JetReconstruction.jl is consistently faster than Fastjet, by 33% for the Durham algorithm and, on average, by 40% for generalised ${k}_{\text{T}}$ $e^{+}e^{-}$  as seen in Table 2.
(a) Durham.
(b) Generalised ${k}_{\text{T}}$ $e^{+}e^{-}$ $p=-1$, $R=1$.
Figure 4: Jet reconstruction times for $e^{+}e^{-}$ algorithms. Reconstruction time for Julia and Fastjet are plotted against the average cluster density of different samples.
Algorithm Runtime Ratio
Durham 1.33
Generalised ${k}_{\text{T}}$ $e^{+}e^{-}$ 1.40
Table 2: Mean ratio of Fastjet reconstruction time to Julia (thus $>1$ indicates Julia is faster) for different $e^{+}e^{-}$ algorithms. The Durham value is averaged over all sample cluster densities. The Generalised ${k}_{\text{T}}$ $e^{+}e^{-}$ is averaged over all sample cluster densities, $p\in(-1,0,1)$ and $R\in(0.2,0.4,1,1.5,2,4)$.
5 Substructure and Taggers
At the LHC jet substructure is a critical component of many physics studies, probing the internal structure of jets, which can be crucial for distinguishing between different types of particles and for identifying new physics signals. In JetReconstruction.jl, since v0.4.4, we have implemented several substructure algorithms and taggers that are commonly used in LHC analyses.
The substructure algorithms soft drop Larkoski_2014 and mass drop Butterworth_2008 are implemented. To use these taggers, we allow users to define a simple structure with the relevant tagging parameters, then to call the appropriate tagger method. The jet filtering Butterworth_2008 and jet trimming Krohn_2010 methods have also been introduced in JetReconstruction.jl. These methods are used to mitigate the effects of pileup and underlying event contamination in jet reconstruction and are used in a similar way to the taggers: a simple struct of parameters is defined and the appropriate method called to obtain a cleaned jet.
The average time to perform the substructure and tagging routines on our benchmark machine is shown in Table 3. Fastjet is faster for filtering and trimming, while JetReconstruction.jl is much faster for the mass drop algorithm.
Method JetReconstruction.jl Fastjet Ratio (Fj/JR)
Jet Filtering 3.37 2.47 0.73
Jet Trimming 2.80 2.26 0.81
Mass Drop 0.26 0.85 3.2
Soft Drop 2.85
Table 3: Time in $\mu\mathrm{s}$ to perform various substructure and tagging operations in JetReconstruction.jl and in Fastjet (jet filtering is with a filter radius of 0.3 and selecting the 3 hardest jets; trimming is uses a trim radius and trim fraction of 0.3, with Cambridge/Aachen reclustering; mass drop uses $\mu=0.67$, $y=0.09$; and soft drop uses $z_{cut}=0.1$, $\beta=2.0$). The average is taken by performing these operations on all inclusive jets above 2GeV in all of our input sample files. Note that soft drop is not available in the core Fastjet release.
6 FCCee Jets and EDM4hep
As an example of how to easily support experiment event data models (EDM) in JetReconstruction.jl we have introduced support for reading EDM4hep Gaede:2022leb events directly. Like visualisation, this support takes advantage of the Julia package manager’s concept of extensions, which means that the additional code to support this is not loaded unless the EDM4hep.jl package is present in the user’s environment. Once this is done it is trivial to call the main jet reconstruction directly using the ReconstructedParticles object from an EDM4hep ROOT file.
On our benchmark machine, using the Durham algorithm, the jet reconstruction can reach 24kHz on a single thread.
7 Conclusions
JetReconstruction.jl has been released as a production package in the Julia ecosystem. It is easy to use and gives better performance than Fastjet in most cases. Feedback from users has been very positive. We intend to continue to develop that package in the future, allowing HEP users to benefit from the advantages of Julia in their analyses that require jet reconstruction.